OpenStudy (anonymous):

2xsquared - 8x - 10 = 0

OpenStudy (aroub):

2x^2-8x-10=0 2(x^2-4x-5)=0 2(x-5)(x+1)=0

OpenStudy (aroub):

x=5 or x=-1

OpenStudy (angela210793):

D=b^2-4ac=64-4*2*(-10)=144 sqr144=12 x1=(-b+sqrD)/2a=8+12)4=5 x2=(-b-sqrD)/2a=(8-12)/4=-1